A wheel starts from rest and rotates with constant angular acceleration to reach an angular speed of 12 rad/s in 3.0 s. Find (a) the magnitude of the angular acceleration of the wheel and (b) the angle in radians through which it rotates in this time interval.
Answer #1

Solution:

Using the rotational dynamics,

a)

= - o/t

= 12 - 0/3

= 4 rad/s

b)

= ot + 1/2t2

= 0 + 0.5*4*32

= 18 radians .

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions